### v1

# Exploring Bias and Prediction Metrics to Characterise the Fairness of Machine Learning for Equity-Centered Public Health Decision-Making: A Narrative Review

Exploring Bias and Prediction Metrics to Characterise the Fairness of Machine Learning for Equity-Centered Public Health Decision-Making: A Narrative Review
1 Supplementary Information
1.1 Glossary of Terms in Public Health, Equity, Fairness, Bias, and Machine Learning
Health
: A state of complete physical, mental, and social well-being and not merely the absence of disease or infirmity. It covers the overall condition of an individual or population, including the presence or absence of illnesses, the level of physical and mental function, and the ability to respond to various factors that affect health.
Public Health
: Refers to the science and art of preventing disease, prolonging life, and promoting health through organized efforts and informed choices of society, organizations, public and private, communities, and individuals.
Equity
: The absence of avoidable or remediable differences among groups of people, whether those groups are defined socially, economically, demographically, or geographically. In health care, equity refers to providing care that does not vary in quality because of personal characteristics such as gender, ethnicity, geographic location, and socio-economic status.
Machine Learning (ML)
: A subset of artificial intelligence that involves the development of algorithms and statistical models that enable computers or programs to perform a task without using explicit instructions, relying instead on patterns and inference derived from data.
Fairness
: In the context of ML and decision-making systems, fairness refers to the attribute that an algorithmic decision or model does not create biased outcomes, ensuring that all individuals or groups are treated equally and justly based on the relevant criteria.
Bias
: Refers to a systematic error or non-random distortion in results or inferences from processes. In data science and machine learning, bias can emerge due to prejudiced assumptions during the model development phase, from the data collection process, or in the algorithmic processing, leading to unfair outcomes.
Algorithmic Bias
: Occurs when a computer program or algorithm reflects the implicit values of the humans who are involved in coding, collecting, selecting, or using data to train the algorithm. This can lead to skewed outputs and discriminatory practices.
Social Determinants of Health (SDOH)
: The conditions in the environments where people are born, live, learn, work, play, worship, and age that affect a wide range of health, functioning, and quality-of-life outcomes and risks.
Health Disparities
: A type of difference in health that is closely linked with social, economic, and/or environmental disadvantage. Health disparities adversely affect groups of people who have systematically experienced greater obstacles to health based on their racial or ethnic group; religion; socio-economic status; gender; age; mental health; cognitive, sensory, or physical disability; sexual orientation or gender identity; geographic location; or other characteristics historically linked to discrimination or exclusion.
Predictive Analytics
: The use of data, statistical algorithms, and machine learning techniques to identify the likelihood of future outcomes based on historical data. In public health, it’s used for predicting disease outbreaks, health trends, and patient outcomes.
Informed Consent
: A process for getting permission before conducting a healthcare intervention on a person. In research, it involves informing participants about the risks, benefits, and potential outcomes of studies or data collection.
Privacy
: In the context of data and digital tools, privacy concerns the right of individuals to control or influence what information related to them may be collected and stored and by whom and to whom that information may be disclosed.
Data Governance
: The overall management of the availability, usability, integrity, and security of the data employed in an organization, with the objective to ensure that data is accurate, available, and accessible.
Disparate Impact
: Refers to practices or policies that may appear neutral but result in a disproportionate impact on a protected group. It’s often used to identify unintentional biases in systems or models.
Predicted Positive Rate (PPR)
: Measures the proportion of individuals in a group predicted to have a positive outcome. It’s useful for identifying bias in predictive models across different demographic groups.
False Discovery Rate (FDR)
: The proportion of false positives within all positive predictions. A high FDR indicates many incorrect positive predictions, suggesting a need for model recalibration.
False Positive Rate (FPR)
: Indicates how frequently healthy individuals are mistakenly identified as diseased. A high FPR can lead to unnecessary stress and testing, highlighting the need for accurate diagnostic criteria.
False Omission Rate (FOR)
: Assesses the rate of false negatives among negative predictions. A high FOR, especially in diagnosing infectious diseases, could lead to untreated spread within communities.
False Negative Rate (FNR)
: Reflects the proportion of actual positives missed by the model. A high FNR, such as in COVID-19 testing, means many infected individuals are wrongly reassured they are virus-free, posing a public health risk.
Group Size Ratio (GSR)
: Examines the representation ratio of different groups within the data. It’s important for ensuring research outcomes are representative and equally applicable across all groups.
Equality of Opportunity
: Ensures individuals from different groups have equal chances of receiving a favorable outcome. It’s crucial in healthcare to ensure decisions are based on medical need rather than the ability to pay.
Equalized Odds
: Aims for similar error rates across groups to ensure fairness in treatment recommendations and other decisions made by predictive models.
Balanced Accuracy
: Balances sensitivity and specificity across groups to ensure that a model accurately identifies needs regardless of demographic factors.
Predictive Parity
: Measures consistency in predictive outcomes across different groups, ensuring that no group is disproportionately flagged by predictive models.
Calibration
: Evaluates the alignment of predicted probabilities with actual outcomes across groups. It’s essential for maintaining the reliability of predictive models in public health and other areas.
Balanced Error Rate (BER)
: Seeks to balance false positives and negatives to prevent bias in predictive modeling, ensuring equitable outcomes across all groups.
Background Negative Subgroup Positive (BNSP) AUC
: Focuses on model accuracy when identifying positive cases against a largely negative backdrop. It’s critical in rare disease identification to avoid overwhelming false alerts.
1.2 Algorithmic Processing
Techniques such as data augmentation and synthetic data generation can improve data diversity [leavy2018gender, zhao2018gender, yang2022algorithmic]. Additionally, methods like re-weighting and re-sampling are critical for achieving fair representation [zelaya2019towards, kamiran2012data, zhang2020fairness].
The model training phase is also susceptible to bias through the selection of model architecture, hyperparameters, and loss functions.
Mitigation strategies include designing models less sensitive to biased features, using fair loss functions, and adjusting training processes to improve fairness [nielsen2020practical, lahoti2020fairness, mandal2020ensuring, kearns2018preventing, agarwal2019fair]. During model calibration, ensuring accurate and fair probabilistic estimates is crucial to avoid biases impacting prediction quality and decision fairness. Approaches for addressing calibration biases involve adjusting output probabilities, equalizing decision thresholds across diverse groups, and properly calibrating model outputs [hardt2016equality, jang2022group, pleiss2017fairness].
1.3 Equitable Predictions in Machine Learning: A Case Study on Diabetes Readmission
1.3.1 Task Setting and Evaluation for Fairness
This case study investigates the fairness of machine learning (ML) model predictions across various demographics—age, gender, and race—using a comprehensive diabetes dataset. Our primary focus is on evaluating the equity of these predictions in determining the risk of 30-day hospital readmissions for diabetic patients. We examine whether ML models offer equitable predictions across different demographic groups, particularly in forecasting 30-day readmissions among diabetes patients. Our objective encompasses not only assessing the accuracy of these models but also scrutinizing their fairness and potential biases to ensure that healthcare predictive analytics equitably benefit all patient groups.
Leveraging data from 130 US hospitals between 1999 and 2008, this study narrows down on 45,715 patient records from an initial pool of 101,766, after filtering out identifiers, features with substantial missing values, and records lacking race or diagnosis information. The final dataset includes 50 attributes, highlighting a class imbalance in the 30-day readmission indicator. Our analysis deploys four ML models—Multi-Layer Perceptron (MLP), Generalized Linear Model (GLM), and Naive Bayes—optimized for accuracy and fairness across demographic lines through specific hyperparameter settings.
The evaluation framework combines accuracy metrics (precision, recall, F1 score, and overall accuracy) with a thorough fairness analysis. By adopting metrics like the disparate impact ratio and predicted positive rates, among others, we aim to provide equitable predictions across all demographics. The study follows a structured evaluation flow: generating model predictions, assessing accuracy, conducting a fairness analysis for sub-groups, examining accuracy within these groups, and iteratively refining models to balance accuracy with fairness.
In a city, over the past year there is a hospital where 10,000 White patients and 5,000 Black patients sought diagnosis for a disease that requires hospitalization. The White population is assumed to be in the majority. We want to examine the privileged and underprivileged groups who get the favourable outcome (hospitalization here), regardless of their population size. The following confusion matrices represent the actual number of patients from each racial subgroup who were diagnosed with the disease and required hospitalization (positive label for ‘hospitalized’ and negative for ‘not hospitalized’).
1.3.2 Results
Table 1: Comparison of Different Models: Multilayer Perceptron (MLP), Generalized Linear Model (GLM), Naive Bayes (NB) Performance Across Precision, Recall, F1 Score, and Accuracy Metrics.
Model Precision Recall F1 Score Accuracy
MLP 0.7978 0.987 0.882 0.819
GLM 0.7929 0.928 0.855 0.8097
Naive Bayes 0.7786 0.823 0.801 0.7404
Table 1 compares three models (MLP, GLM, Naive Bayes) on metrics like Precision, Recall, F1 Score, and Accuracy. We observe that MLP performs best across all metrics. It indicates that MLP is most effective in balancing correct predictions and minimizing false positives and negatives. GLM follows closely, while Naive Bayes lags behind, possibly due to its assumption of feature independence. The choice of model should also consider factors like computational cost and interpretability beyond these metrics.
While the performance metrics of these models are key to measure their effectiveness, assessing their impact across various demographic groups is vital for ensuring fairness. Fairness entails that these models’ predictions do not systematically advantage or disadvantage any group based on sensitive attributes like age, gender, race, or socio-economic status. Evaluating model performance with a lens on demographic equity is essential to uncovering biases that could result in unfair outcomes, thus ensuring the models operate equitably without reinforcing or worsening existing disparities.
Table 2: Evaluation of Fairness Metrics (Predicted Positive Group Rate (PPGR), False Positive Rate (FPR), False Omission Rate (FOR), False Negative Rate (FNR)) Across Gender (Female and Male) Groups for MLP, GLM, and NB Models. Bold indicates the best score within each gender group, with lower error rates (FPR, FOR, FNR) preferred and higher PPGR values suggesting more equitable positive predictions.
Gender Model PPGR FPR FOR FNR
Female MLP 0.37 0.16 0.25 0.31
GLM 0.38 0.17 0.27 0.32
NB 0.35 0.13 0.35 0.45
Male MLP 0.38 0.07 0.23 0.27
GLM 0.42 0.08 0.24 0.29
NB 0.4 0.12 0.31 0.42
Table 2 presents a nuanced view of fairness metrics across gender groups for different models, revealing disparities in model biases. For females, the MLP model demonstrates a balance in predicted PPGR and maintains lower error rates compared to GLM and NB, indicating a slightly less biased performance towards female subjects. However, for males, the MLP model shows superior fairness metrics, with the highest PPGR and the lowest rates of FPR, FOR, and FNR, suggesting it is the least biased model across both genders. Conversely, the NB model exhibits higher error rates for both genders, particularly in the FOR and FNR for females, and overall for males, pointing to a greater bias. This analysis indicates that while no model is entirely free of bias, MLP appears to be the most equitable, whereas NB could be considered the most biased, with GLM occupying a middle ground.
Table 3: Comparative Analysis of Fairness Metrics by Race for MLP, GLM, and NB Models. The table highlights disparities in model performance across racial groups, focusing on fairness through PPGR, FPR, FNR and FOR metrics. Bold figures indicate the most favorable outcomes, with lower error rates (FPR, FOR, FNR) preferred and higher PPGR values suggesting more equitable positive predictions
Race Model PPGR FPR FOR FNR
African American MLP 0.38 0.08 0.24 0.29
GLM 0.39 0.08 0.24 0.29
NB 0.36 0.14 0.31 0.40
Asian MLP 0.56 0.01 0.23 0.17
GLM 0.41 0.05 0.27 0.29
NB 0.36 0.04 0.48 0.46
Caucasian MLP 0.37 0.06 0.24 0.31
GLM 0.39 0.07 0.24 0.29
NB 0.36 0.13 0.31 0.41
Hispanic MLP 0.38 0.11 0.35 0.39
GLM 0.39 0.08 0.31 0.35
NB 0.34 0.11 0.39 0.47
Other MLP 0.5 0.11 0.34 0.27
GLM 0.41 0.13 0.30 0.34
NB 0.45 0.11 0.40 0.35
Table 3 disparities in model performance across racial groups, focusing on fairness metrics. Asians benefit most notably from the MLP model’s precision, experiencing the most fir treatment with the lowest FPR and highest PPGR, highlighting its superior predictive equity. In contrast, the Hispanic and Asian groups face the greatest challenges with the NB model, which shows a tendency to underpredict positive outcomes for these demographics, indicating significant biases.
Overall, in these results, we observe that the MLP model emerges as the most equitable, especially for Asians, by achieving an optimal balance between fairness and accuracy across racial groups. The GLM model displays consistent fairness metrics, offering a stable, albeit less exceptional, performance that does not heavily favor or disadvantage any particular group. On the other hand, the NB model requires refinement to address its bias, particularly in improving fairness for Hispanic and Asian populations, where it currently lags in predictive accuracy and equity.